# A geometric approach toconjugation-invariant random permutations

### v3

@@ 2.2 Robinson–Schensted shape
Let $\tau$ be a permutation of $[n]$. If $x\geq 0$, we call $x$-decrasing subsequence any union of $\lfloor x\rfloor$ individually decreasing subsequences, and we define $\mathrm{LDS}_{x}(\tau)$ as the maximal size of an $x$-decreasing subsequence of $\tau$.
The study of $x$-decreasing subsequences in permutations is partly motivated by their fundamental link with the Robinson–Schensted correspondence. This is a one-to-one correspondence between permutations and pairs of standard Young tableaux with the same shape. This shape, which is a Young diagram, is called the RS shape of the permutation: it appears in several domains, such as integrable probability or representation theory. A well-known theorem of Greene [Gre74] states that for any integer $k$, the number of boxes in the first $k$ columns of the RS shape of $\tau$ equals $\mathrm{LDS}_{k}(\tau)$. Equivalently, the length of the $k$-th column equals $\mathrm{LDS}_{k}(\tau)-\mathrm{LDS}_{k\!-\!1}(\tau)$. A similar connection can be made between increasing subsequences of $\tau$ and the row lengths of its RS shape.
@@
A classical property of the Robinson–Schensted correspondence is that it induces a bijection between involutions and (single) standard Young tableaux. If $(\tau_{n})$ is a sequence of uniformly random involutions, the proofs of [LS77] and [VK77] can then be adapted to show that Theorem 2.6 still holds (see e.g. Equation (24) in [Mel11] and the discussion below). More generally if $(\tau_{n})$ is a sequence of conjugation-invariant random permutations, Kammoun proved that Theorem 2.6 still holds under the assumption that the number of cycles in $\tau_{n}$ is sublinear [Kam18, Theorem 1.8]. He later conjectured that this assumption could be lifted, provided we take into account the proportion of fixed points [Kam23, Conjectures 6 and 7]. The main difficulty is that none of the previous approaches (via hook-length formula [LS77], representation theory [IO02, Mel11], coupling [Kam18]…) work in this wider setting. Here, using our geometric construction, we solve this conjecture.
@@
Let us briefly discuss the strategy of proof. As in Section 2.1, $k$-decreasing subsequences of $\tau={\rm Perm}\left(\mathcal{P}\right)$ have a nice visual interpretation: they correspond to unions of $k$ down-right paths in $\mathcal{P}$. In the continuity of [DZ95], where the longest monotone subsequences of locally uniform permutations were studied with geometric tools, Sjöstrand [Sjö23] recently studied the limit RS shape of locally uniform permutations using more advanced analysis. His approach broadly consisted in computing the “score” associated with a bundle of decreasing curves, by decomposing the unit square into small regions where the sampling density is almost constant and the curves are almost parallel. In 3.3 we show that the points of our geometric construction, when restricted to any rectangle outside the diagonal of the unit square, are i.i.d. uniform. This simple fact is at the basis of how we apply and adapt the results and methods of [Sjö23].
The results of this section are proved in Section 5.
@@ 2.3 Records
Let $\tau$ be a permutation of $[n]$. We say that a position $i\in[n]$ is a high record (or left-to-right maximum) for $\tau$ if for any $j<i$, we have $\tau(j)<\tau(i)$. We say that it is a low record (or left-to-right minimum) for $\tau$ if for any $j<i$, we have $\tau(j)>\tau(i)$. We denote by $\mathrm{rec}^{\uparrow}\left(\tau\right)$ and $\mathrm{rec}^{\downarrow}\left(\tau\right)$ the numbers of high and low recods in $\tau$, respectively. Records and their many variants are standard statistics in enumerative combinatorics.
The numbers of low and high records in uniformly random permutations are well understood: thanks to a classical bijection they follow the same law as the number of cycles (see e.g. [ABNP16, Section 1.2]), and this yields the following asymptotic normality (found e.g. in [ABT03, Equation (1.31)]).
@@
Records can be interpreted nicely when the permutation is obtained from a family of points in the plane. If $\tau={\rm Perm}\left(\mathcal{P}\right)$, a point $Z=(X,Y)\in\mathcal{P}$ corresponds to a high record in $\tau$ when no other point lies in its up-left corner $[0,X]\times[Y,1]$. Similarly, $Z$ corresponds to a low record in $\tau$ when no other point lies in its down-left corner $[0,X]\times[0,Y]$. We write $\mathrm{rec}^{\uparrow}\left(\mathcal{P}\right)$ and $\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)$ as shortcuts for $\mathrm{rec}^{\uparrow}\left({\rm Perm}\left(\mathcal{P}\right)\right)$ and $\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\mathcal{P}\right)\right)$.
@@
It would be interesting to find the interpolation between the two items of 2.12. However this would notably require understanding the interaction between $2$-cycles and other cycles, and its effect on the law of low records, which seems difficult to grasp.
@@
The results of this section are proved in Section 6.
@@ 2.4 Pattern counts
[…]
Pattern counts in uniformly random permutations are known to be asymptotically normal since the works of Bóna for monotone patterns [Bón10], and later Janson et al. for all patterns:
@@
In [Fér13], Féray established a central limit theorem for pattern counts in Ewens random permutations, and conjectured its non-degeneracy. Kammoun [Kam22, Proposition 31] and Hamaker and Rhoades [HR22, Theorem 8.8] could generalize the non-degenerate asymptotic normality of Theorem 2.13 to conjugation-invariant permutations, under certain conditions on their cycles. It was conjectured in [HR22, Problem 9.9] that those conditions could be lifted, and that the asymptotic variance would only depend on the proportions of fixed points and $2$-cycles.
Here, using the geometric construction of 1.1, we extend the convergence of Theorem 2.13 to most conjugation-invariant permutations. If $\tau_{n}$ is a uniform $t$-cyclic permutation then we can write, for any $\pi\in\mathfrak{S}_{r}$:
@@
with the notation of Section 1.2. Thanks to the dependency graph $\mathcal{L}_{t}$ defined in Section 1.3, we derive a limit theorem for $X_{\pi}(\tau_{n})$. More precisely, Stein’s method yields explicit bounds on the speed of convergence. In the following, we denote by $d_{K}$ the Kolmogorov distance between two probability distributions.
@@
Furthermore if $\mathrm{Var}\left[X_{\pi}(\tau_{n})\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ for some $\sigma\geq 0$ as $n\to\infty$, then
@@
Thanks to [HR22, Proposition 7.2 and Theorem 8.14], the condition $\mathrm{Var}\left[X_{\pi}(\tau_{n})\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ holds under the simple assumption that $t_{1}^{(n)}/n$ and $t_{2}^{(n)}/n$ converge. We may still want to find an explicit formula for $\sigma$, and to understand in which cases $\sigma>0$. For this we use the method of $U$-statistics developed by Hoeffding [Hoe48], as was done in [JNZ15]. Before stating our results, we need a few definitions.
Fix $r\in\mathbb{N}^{*}$ and $p_{1}\in[0,1]$. Then let $\hat{Z}_{1},\dots,\hat{Z}_{r}$ be i.i.d. variables distributed under $p_{1}\mathrm{Leb}_{\Delta}+(1\!-\!p_{1})\mathrm{Leb}_{[0,1]^{2}}$, where $\mathrm{Leb}_{\Delta}$ denotes the Lebesgue measure on the diagonal $\Delta$ of $[0,1]^{2}$. Define, for any $z\in[0,1]^{2}$:
@@
Let $(\tau_{n})_{n}$ be a sequence of conjugation-invariant random permutations, where $\tau_{n}$ has size $n$ and (random) cycle type $t^{(n)}$. Suppose that $t_{1}^{(n)}=np_{1}+o_{\mathbf{P}}(\sqrt{n})$ and $2t_{2}^{(n)}=np_{2}+o_{\mathbf{P}}(n)$ as $n\to\infty$, for some $p_{1},p_{2}\in[0,1]$. Then the convergence of Theorem 2.15 holds.
Note that Theorem 2.15 states a joint convergence, rather than the marginal convergence stated in Theorem 2.14. We stress that the true novelty of Theorem 2.15 lies in the computation of the variance, carried out by the method of $U$-statistics, and not in the joint convergence. Indeed, Theorem 2.14 could have been stated for joint convergence, but we chose not to do so for convenience.
@@
If $p_{1}<1$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$ satifies $\pi=\pi^{-1}$, then $\Sigma_{\pi,\pi}^{p_{1},p_{2}}>0$.
When fixed points vanish, i.e. when $p_{1}=0$, Theorem 2.15 greatly simplifies. This allows us to prove non-degeneracy for any non-trivial pattern and to compute the dimension spanned by the limiting gaussian variable, as in Theorem 2.13 of [JNZ15]. We use the notation of Theorem 2.15 and drop the index $p_{1}$ when it is null.
@@
The matrix $\Sigma^{p_{2}}$ has rank $(r\!-\!1)^{2}$ for any $p_{2}<1$, and rank $r(r\!-\!1)/2$ for $p_{2}=1$. Moreover $\Sigma_{\pi,\pi}^{p_{2}}>0$ for any $p_{2}\in[0,1]$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$.
@@
In parallel to this work, Féray and Kammoun [FK23] were able to prove a more general result than Theorem 2.15 and 2.17, using the method of weighted dependency graphs. They could prove non-degeneracy for any (classical) non-trivial pattern, i.e. that $\Sigma_{\pi,\pi}^{p_{1},p_{2}}>0$ for any $p_{1}<1$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$. However, bounds on the speed of convergence as in our Theorem 2.14 seem to be out of reach for their methods.
@@
Our method of (non-weighted) dependency graphs also allows the application of [Jan04, Corollary 2.2] to Eq. (1) for large deviation estimates. Namely, for any $t>0$:
$\mathbf{P}\left(\big{.}\left|X_{\pi}(\tau_{n})-\mathbb{E}\left[X_{\pi}(\tau_{n})\right]\right|\geq t\right)\leq 2\exp\left(\frac{-2(r-1)!^{2}}{3}\frac{t^{2}}{n^{2r-1}}\right)$
@@
A popular generalization of patterns is that of vincular patterns, where we require some positions of the pattern to be adjacent. Descents, peaks and valleys are classical examples of vincular patterns. Asymptotic normality of vincular pattern counts in uniform permutations was established in [Hof18]. Regarding conjugation-invariant permutations, the convergence results of [Kam22, HR22, FK23] all hold within the general context of vincular patterns. Unfortunately, it seems unlikely that our methods would work for vincular patterns.
The results of this section are proved in Section 7.
@@ 3 Properties of the geometric construction
In this section, we present several interesting facts about the geometric construction of 1.1. These will be essential to prove the results of this paper, generally by reducing the study of conjugation-invariant permutations to uniformly random permutations. We use the notation of Sections 1.2 and 1.3.
@@
where each $\mathcal{P}^{(i)}$ is a family of i.i.d. uniform points in $[0,1]^{2}$, of (deterministic) sizes $n_{1},n_{2},n_{3}$ bounded between $n/3\!-\!1$ and $n/3\!+\!1$.
@@
  • For each $p$ such that $p=1\!\mod 3$ and $1\leq k\leq t_{p}$, the points of $\mathcal{P}_{p,k}$ can be put in the three subsets by simply alternating, but then one subset will get a surplus. Thus two of the subsets $\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$ have size $(p\!-\!1)/3$, the other one has size $31+(p\!-\!1)/$ 3, and they each contain i.i.d. uniform points. Namely if $(i_{1},i_{2},i_{3})$ is a permutation of $(1,2,3)$ then the choice
@@
    $\bigcup_{p=1\mod 3,1\leq k\leq t_{p}}\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$
@@
    $\bigcup_{p=2\mod 3,1\leq k\leq t_{p}}\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$
@@
Thanks to 3.1, we can apply Chernoff-type bounds to control the number of points in any zone of the plane. Recall that $\check{\mathcal{P}}:=\mathcal{P}\setminus\Delta$ is a geometric construction for $\check{t}:=(0,t_{2},\dots)$ of size $\check{n}:=n-t_{1}$.
@@
Let $t$ be a cycle type of size $n$ and $\mathcal{P}$ be a geometric construction. If $C$ is a measurable subset of $[0,1]^{2}$ then $\left|\check{\mathcal{P}}\cap C\right|$ is distributed like a (correlated) sum of ${\rm Binom}\left(n_{i},\mathrm{Leb}(C)\right)$ variables, where $\left|n_{i}-\check{n}/3\right|\leq 1$ for $i=1,2,3$. In particular for any $\delta>0$ there exists a constant $c>0$, which depends only on $\delta$ and $\mathrm{Leb}(C)$, such that:
@@
It suffices to prove the lemma for each set $\mathcal{P}_{p,k}$. Fix $(p,k)$. Note that if some point $\big{(}U_{p,k}^{l},U_{p,k}^{l+1}\big{)}$ is in $I\times J$ then a.s. neither $\big{(}U_{p,k}^{l-1},U_{p,k}^{l}\big{)}$ nor $\big{(}U_{p,k}^{l+1},U_{p,k}^{l+2}\big{)}$ are in $I\times J$, where all exponents are taken modulo $p$ in $[p]$. Moreover $\big{(}U_{p,k}^{l},U_{p,k}^{l+1}\big{)}$ is independent from
@@
{.$\left\{\Bigg{.}\mathcal{P}_{p,k}\cap(I\times J)=\left\{\big{(}U_{p,k}^{l_{1}},U_{p,k}^{l_{1}+1}\big{)},\dots,\big{(}U_{p,k}^{l_{j}},U_{p,k}^{l_{j}+1}\big{)}\right\}\right\}$
@@
The conditional law of $\mathcal{P}$ given $\{U_{i}\,:\,i\in\mathcal{V}_{t}\}$ can be a.s. re-described as follows. Sort $\{U_{i}\,:\,i\in\mathcal{V}_{t}\}$ as $U_{(1)}<\dots<U_{(n)}$, let $\rho$ be a uniformly random bijection from $[n]$ to $\mathcal{V}_{t}$, then set $V_{i}:=U_{(\rho^{-1}(i))}$ for any $i\in\mathcal{V}_{t}$, and finally $\mathcal{P}=\left\{\big{.}(V_{i},V_{\mathfrak{s}(i)})\,:\,i\in\mathcal{V}_{t}\right\}$. The first claim then follows from the same argument as in the proof of 1.1.
For the second claim, notice that $\mathcal{P}\cap(I\times J)$ is a random subset of the finite grid $\left\{\left(U_{(i)},U_{(j)}\right):i,j\in[n]\right\}\cap(I\times J)$. Consider $0\leq m\leq n$ and indices $i_{1},\dots,i_{m},j_{1},\dots,j_{m}\in[n]$ such that for each $r\in[m]$, $U_{(i_{r})}\in I$ and $U_{(j_{r})}\in J$. The sets of indices $\{i_{1},\dots,i_{m}\}$ and $\{j_{1},\dots,j_{m}\}$ are then a.s. disjoint. Furthermore the probability

### v4

@@ 2.2 Robinson–Schensted shape
Let $\tau$ be a permutation of $[n]$. If $x\geq 0$, we call $x$-decreasing subsequence any union of $\lfloor x\rfloor$ individually decreasing subsequences, and we define $\mathrm{LDS}_{x}(\tau)$ as the maximal size of an $x$-decreasing subsequence of $\tau$.
The study of $x$-decreasing subsequences in permutations is partly motivated by their fundamental link with the Robinson–Schensted correspondence. This is a one-to-one correspondence between permutations and pairs of standard Young tableaux with the same shape. This shape, which is a Young diagram, is called the RS shape of the permutation: it appears in several domains, such as integrable probability or representation theory. A well-known theorem of Greene [G74] states that for any integer $k$, the number of boxes in the first $k$ columns of the RS shape of $\tau$ equals $\mathrm{LDS}_{k}(\tau)$. Equivalently, the length of the $k$-th column equals $\mathrm{LDS}_{k}(\tau)-\mathrm{LDS}_{k\!-\!1}(\tau)$. A similar connection can be made between increasing subsequences of $\tau$ and the row lengths of its RS shape.
@@
A classical property of the Robinson–Schensted correspondence is that it induces a bijection between involutions and (single) standard Young tableaux. If $(\tau_{n})$ is a sequence of uniformly random involutions, the proofs of [LS77] and [VK77] can then be adapted to show that Theorem˜2.6 still holds (see e.g. Equation (24) in [M11] and the discussion below). More generally if $(\tau_{n})$ is a sequence of conjugation-invariant random permutations, Kammoun proved that Theorem˜2.6 still holds under the assumption that the number of cycles in $\tau_{n}$ is sublinear [K18, Theorem 1.8]. He later conjectured that this assumption could be lifted, provided we take into account the proportion of fixed points [K23, Conjectures 6 and 7]. The main difficulty is that none of the previous approaches (via hook-length formula [LS77], representation theory [IO02, M11], coupling [K18]…) work in this wider setting. Here, using our geometric construction, we solve this conjecture.
@@
Let us briefly discuss the strategy of proof. As in Section˜2.1, $k$-decreasing subsequences of $\tau={\rm Perm}\left(\mathcal{P}\right)$ have a nice visual interpretation: they correspond to unions of $k$ down-right paths in $\mathcal{P}$. In the continuity of [DZ95], where the longest monotone subsequences of locally uniform permutations were studied with geometric tools, Sjöstrand [S23] recently studied the limit RS shape of locally uniform permutations using more advanced analysis. His approach broadly consisted in computing the “score” associated with a bundle of decreasing curves, by decomposing the unit square into small regions where the sampling density is almost constant and the curves are almost parallel. In ˜3.3 we show that the points of our geometric construction, when restricted to any rectangle outside the diagonal of the unit square, are i.i.d. uniform. This simple fact is at the basis of how we apply and adapt the results and methods of [S23].
The results of this section are proved in Section˜5.
@@ 2.3 Records
Let $\tau$ be a permutation of $[n]$. We say that a position $i\in[n]$ is a high record (or left-to-right maximum) for $\tau$ if for any $j<i$, we have $\tau(j)<\tau(i)$. We say that it is a low record (or left-to-right minimum) for $\tau$ if for any $j<i$, we have $\tau(j)>\tau(i)$. We denote by $\mathrm{rec}^{\uparrow}\left(\tau\right)$ and $\mathrm{rec}^{\downarrow}\left(\tau\right)$ the numbers of high and low records in $\tau$, respectively. Records and their many variants are standard statistics in enumerative combinatorics.
The numbers of low and high records in uniformly random permutations are well-understood: thanks to a classical bijection they follow the same law as the number of cycles (see e.g. [ABNP16, Section 1.2]), and this yields the following asymptotic normality (found e.g. in [ABT03, Equation (1.31)]).
@@
Records can be interpreted nicely when the permutation is obtained from a family of points in the plane. If $\tau={\rm Perm}\left(\mathcal{P}\right)$, a point $Z=(X,Y)\in\mathcal{P}$ corresponds to a high record in $\tau$ when no other point lies in its up-left corner $0,X]\times[Y,1]$. Similarly, $Z$ corresponds to a low record in $\tau$ when no other point lies in its down-left corner $0,X]\times[0,Y]$. We write $\mathrm{rec}^{\uparrow}\left(\mathcal{P}\right)$ and $\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)$ as shortcuts for $\mathrm{rec}^{\uparrow}\left({\rm Perm}\left(\mathcal{P}\right)\right)$ and $\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\mathcal{P}\right)\right)$.
@@
It would be interesting to find the interpolation between the two items of ˜2.12. However this would notably require understanding the interaction between $2$-cycles and other cycles, and its effect on the law of low records, which seems difficult to grasp.
@@
The results of this section are proved in Section˜6.
@@ 2.4 Pattern counts
where $\binom{[n]}{r}$ denotes the subsets of $[n]$ of size $r$. For example, $X_{2\,1}(\tau)$ counts the number of inversions in $\tau$, and $X_{1\,2\,\dots\,r}(\tau)$ counts the number of increasing subsequences of length $r$ in $\tau$.
Pattern counts in uniformly random permutations are known to be asymptotically normal since the works of Bóna for monotone patterns [B10], and later Janson et al. for all patterns:
@@
In [F13], Féray established a central limit theorem for pattern counts in Ewens random permutations, and conjectured its non-degeneracy. Kammoun [K22, Proposition 31] and Hamaker and Rhoades [HR22, Theorem 8.8] could generalize the non-degenerate asymptotic normality of Theorem˜2.13 to conjugation-invariant permutations, under certain conditions on their cycles. It was conjectured in [HR22, Problem 9.9] that those conditions could be lifted, and that the asymptotic variance would only depend on the proportions of fixed points and $2$-cycles.
Here, using the geometric construction of ˜1.3, we extend the convergence of Theorem˜2.13 to most conjugation-invariant permutations. If $\tau_{n}$ is a uniform $t$-cyclic permutation then we can write, for any $\pi\in\mathfrak{S}_{r}$:
@@
with the notation of Section˜1.2. Thanks to the dependency graph $\mathcal{L}_{t}$ defined in Section˜1.3, we derive a limit theorem for $X_{\pi}(\tau_{n})$. More precisely, Stein’s method yields explicit bounds on the speed of convergence. In the following, we denote by $d_{K}$ the Kolmogorov distance between two probability distributions.
@@
Furthermore, if $\mathrm{Var}\left[X_{\pi}(\tau_{n})\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ for some $\sigma\geq 0$ as $n\to\infty$, then
@@
Thanks to [HR22, Proposition 7.2 and Theorem 8.14], the condition $\mathrm{Var}\left[X_{\pi}(\tau_{n})\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ holds under the simple assumption that $t_{1}^{(n)}/n$ and $t_{2}^{(n)}/n$ converge. We may still want to find an explicit formula for $\sigma$, and to understand in which cases $\sigma>0$. For this we use the method of $U$-statistics developed by Hoeffding [H48], as was done in [JNZ15]. Before stating our results, we need a few definitions.
Fix $r\in\mathbb{N}^{*}$ and $p_{1}\in[0,1]$. Then let $\hat{Z}_{1},\dots,\hat{Z}_{r}$ be i.i.d. variables distributed under $p_{1}\mathrm{Leb}_{\Delta}+(1\!-\!p_{1})\mathrm{Leb}_{[0,1]^{2}}$, where $\mathrm{Leb}_{\Delta}$ denotes the Lebesgue measure on the diagonal $\Delta$ of $0,1]^{2}$. Define, for any $z\in[0,1]^{2}$:
@@
Let $(\tau_{n})_{n}$ be a sequence of conjugation-invariant random permutations, where $\tau_{n}$ has size $n$ and (random) cycle type $t^{(n)}$. Suppose that $t_{1}^{(n)}=np_{1}+o_{\mathbf{P}}(\sqrt{n})$ and $2t_{2}^{(n)}=np_{2}+o_{\mathbf{P}}(n)$ as $n\to\infty$, for some $p_{1},p_{2}\in[0,1]$. Then the convergence of Theorem˜2.15 holds.
Note that Theorem˜2.15 states a joint convergence, rather than the marginal convergence stated in Theorem˜2.14. We stress that the true novelty of Theorem˜2.15 lies in the computation of the variance, carried out by the method of $U$-statistics, and not in the joint convergence. Indeed, Theorem˜2.14 could have been stated for joint convergence, but we chose not to do so for convenience.
@@
If $p_{1}<1$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$ satisfies $\pi=\pi^{-1}$, then $\Sigma_{\pi,\pi}^{p_{1},p_{2}}>0$.
When fixed points vanish, i.e. when $p_{1}=0$, Theorem˜2.15 greatly simplifies. This allows us to prove non-degeneracy for any non-trivial pattern and to compute the dimension spanned by the limiting Gaussian variable, as in Theorem˜2.13 of [JNZ15]. We use the notation of Theorem˜2.15 and drop the index $p_{1}$ when it is null.
@@
The matrix $\Sigma^{p_{2}}$ has rank $(r\!-\!1)^{2}$ for any $p_{2}<1$, and rank r(r1)/2r(r\!-\!1)/2 for $p_{2}=1$. Moreover, $\Sigma_{\pi,\pi}^{p_{2}}>0$ for any $p_{2}\in[0,1]$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$.
@@
In parallel to this work, Féray and Kammoun [FK23] were able to prove a more general result than Theorem˜2.15 and ˜2.17, using the method of weighted dependency graphs. They could prove non-degeneracy for any (classical) non-trivial pattern, i.e. that $\Sigma_{\pi,\pi}^{p_{1},p_{2}}>0$ for any $p_{1}<1$ and $\pi\in\mathfrak{S}_{r}$, $r\geq 2$. However, bounds on the speed of convergence as in our Theorem˜2.14 seem to be out of reach for their methods.
@@
Our method of (non-weighted) dependency graphs also allows the application of [J04, Theorem 2.1] to Equation˜1 for large deviation estimates. Namely, for any $t>0$:
$\mathbf{P}\left(\big.\left|X_{\pi}(\tau_{n})-\mathbb{E}\left[X_{\pi}(\tau_{n})\right]\right|\geq t\right)\leq 2\exp\left(\frac{-2(r-1)!}{3}\frac{t^{2}}{n^{2r-1}}\right)$
@@
A popular generalization of patterns is that of vincular patterns, where we require some positions of the pattern to be adjacent. Descents, peaks and valleys are classical examples of vincular patterns. Asymptotic normality of vincular pattern counts in uniform permutations was established in [H18]. Regarding conjugation-invariant permutations, the convergence results of [FK23, HR22, K22] all hold within the general context of vincular patterns. Unfortunately, it seems unlikely that our methods would work for vincular patterns.
The results of this section are proved in Section˜7.
@@ 3 Properties of the geometric construction
In this section, we present several interesting facts about the geometric construction of ˜1.3. These will be essential to prove the results of this paper, generally by reducing the study of conjugation-invariant permutations to uniformly random permutations. We use the notation of Sections˜1.2 and 1.3.
@@
where each $\mathcal{P}^{(i)}$ is a family of i.i.d. uniform points in $0,1]^{2}$, of (deterministic) sizes $n_{1},n_{2},n_{3}$ bounded between n/31n/3\!-\!1 and n/3+1n/3\!+\!1.
@@
  • For each $p$ such that $p=1\!\mod 3$ and $1\leq k\leq t_{p}$, the points of $\mathcal{P}_{p,k}$ can be put in the three subsets by simply alternating, but then one subset will get a surplus. Thus two of the subsets $\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$ have size (p1)/3(p\!-\!1)/3, the other one has size 1+(p1)/31+(p\!-\!1)/3, and they each contain i.i.d. uniform points. Namely if $(i_{1},i_{2},i_{3})$ is a permutation of $(1,2,3)$ then the choice
@@
    $\bigcup_{p=1\hskip-5.7361pt\mod 3,1\leq k\leq t_{p}}\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$
@@
    $\bigcup_{p=2\hskip-5.7361pt\mod 3,1\leq k\leq t_{p}}\mathcal{P}_{p,k}\cap\mathcal{P}^{(i)}$
@@
Thanks to ˜3.1, we can apply Chernoff-type bounds to control the number of points in any zone of the plane. Recall that $\check{\mathcal{P}}:=\mathcal{P}\setminus\Delta$ is a geometric construction for $\check{t}:=(0,t_{2},\dots)$ of size $\check{n}:=n-t_{1}$.
@@
Let $t$ be a cycle type of size $n$ and $\mathcal{P}$ be a geometric construction. If $C$ is a measurable subset of $0,1]^{2}$ then $\left|\check{\mathcal{P}}\cap C\right|$ is distributed like a (correlated) sum of ${\rm Binom}\left(n_{i},\mathrm{Leb}(C)\right)$ variables, where $\left|n_{i}-\check{n}/3\right|\leq 1$ for $i=1,2,3$. In particular for any $\delta>0$ there exists a constant $c>0$, which depends only on $\delta$ and $\mathrm{Leb}(C)$, such that:
@@
It suffices to prove the lemma for each set $\mathcal{P}_{p,k}$. Fix $(p,k)$. Note that if some point $\big(U_{p,k}^{l},U_{p,k}^{l+1}\big)$ is in $I\times J$ then a.s. neither $\big(U_{p,k}^{l-1},U_{p,k}^{l}\big)$ nor $\big(U_{p,k}^{l+1},U_{p,k}^{l+2}\big)$ are in $I\times J$, where all exponents are taken modulo $p$ in $[p]$. Moreover, $\big(U_{p,k}^{l},U_{p,k}^{l+1}\big)$ is independent of
@@
$\left\{\Bigg.\mathcal{P}_{p,k}\cap(I\times J)=\left\{\big(U_{p,k}^{l_{1}},U_{p,k}^{l_{1}+1}\big),\dots,\big(U_{p,k}^{l_{j}},U_{p,k}^{l_{j}+1}\big)\right\}\right\}$
@@
The conditional law of $\mathcal{P}$ given $\{U_{i}\,:\,i\in\mathcal{V}_{t}\}$ can be a.s. re-described as follows. Sort $\{U_{i}\,:\,i\in\mathcal{V}_{t}\}$ as $U_{(1)}<\dots<U_{(n)}$, let $\rho$ be a uniformly random bijection from $[n]$ to $\mathcal{V}_{t}$, then set $V_{i}:=U_{(\rho^{-1}(i))}$ for any $i\in\mathcal{V}_{t}$, and finally $\mathcal{P}=\left\{\big.(V_{i},V_{\mathfrak{s}(i)})\,:\,i\in\mathcal{V}_{t}\right\}$. The first claim then follows from the same argument as in the proof of ˜1.3.
For the second claim, notice that $\mathcal{P}\cap(I\times J)$ is a random subset of the finite grid $\left\{\left(U_{(i)},U_{(j)}\right):i,j\in[n]\right\}\cap(I\times J)$. Consider $0\leq m\leq n$ and indices $i_{1},\dots,i_{m},j_{1},\dots,j_{m}\in[n]$ such that for each $r\in[m]$, $U_{(i_{r})}\in I$ and $U_{(j_{r})}\in J$. The sets of indices $\{i_{1},\dots,i_{m}\}$ and $\{j_{1},\dots,j_{m}\}$ are then a.s. disjoint. Furthermore, the probability